Problem Page A jet travels 6096 against a jetstream in 8 hours and 7056 with the jetstream in the same amount of time. What is the rate of the jet in still air and what is the rate of the jetstream?
Answer by ewatrrr(24785)


Hi
D = rt 0r r = D/t
(r+c) = 7056/8 = 882
(r-c) = 6096/8 = 762
2r = 1644
r = 822mph, jet's speed jet in still air
rate of the Jetstream = 60mph 882-822